Required information

[The following information applies to the questions displayed below.]

Alexa owns a condominium near Cocoa Beach in Florida. This year, she incurs the following expenses in connection with her condo:

  

Insurance $ 2,850
Mortgage interest 7,775
Property taxes 2,680
Repairs & maintenance 1,825
Utilities 4,200
Depreciation 16,625

During the year, Alexa rented out the condo for 100 days. She did not use the condo at all for personal purposes during the year. Alexa’s AGI from all sources other than the rental property is $200,000. Unless otherwise specified, Alexa has no sources of passive income.

Assuming Alexa receives $23,400 in gross rental receipts, answer the following questions: (Leave no answer blank. Enter zero if applicable.)

a. What effect does the rental activity have on her AGI for the year?

b. Assuming that Alexa’s AGI from other sources is $90,000, what effect does the rental activity have on Alexa’s AGI? Alexa makes all decisions with respect to the property
Answer #1

Answer (a):

Total expenses = 2850 + 7775 + 2680 + 1825 + 4200 + 16625 = $35,955

Assuming Alexa receives $23,400 in gross rental receipts, she will incur losses = 23400 -35955 = ($12,555)

Alexa’s AGI from all sources other than the rental property is $200,000.

As her AGI from all sources other than the rental property is higher than $150,000, she will not be able to deduct any amount of loss against portfolio or active income.

Hence the loss is not deductible and rental activity has no effect on her AGI.

Answer (b):

Assuming that Alexa’s AGI from other sources is $90,000.

Since her AGI is less $100,000, she will be able to deduct $12,555

Hence rental activity loss will reduce her AGI to = 90000 - 12555 = $77,445

The rental activity loss will reduce her AGI by $12,555 and her AGI would be  $77,445